# Slow Inter-area Electro-mechanical Oscillations Revisited: Structural Property of Complex Multi-area Electric Power Systems

### v1

@@ Abstract
This paper introduces a physically-intuitive notion of inter-area dynamics in systems comprising multiple interconnected energy conversion modules. The ideas build on an earlier general approach to setting their structural properties by modeling first stand-alone modular dynamics starting from the fundamental relations between energy stored in modules (components, areas), and constraining explicitly their Tellegen’s quantities, power and rate of change of power, in particular. In this paper we derive, by following the same principles, a transformed state-space model for a general nonlinear system. Using this model we show the existence of an area-level interaction variable, intVar, whose rate of change depends solely on the area internal power imbalance. Given these structural properties of stand-alone modules, we define in this paper for the first time an inter-area variable as the difference of power wave incident to tie-line from Area I and the power reflected into tie-lie from Area II. Notably, these power waves represent the rate of change of intVars associated with the two interconnected areas. We illustrate these notions using a linearized case of two lossless inter-connected areas, and show the existence of a new inter-area mode when the areas get connected. We suggest that lessons learned in this paper open possibilities for computationally-efficient modeling and control of inter-area oscillations, and offer further the basis for modeling and control of dynamics in changing systems comprising faster energy conversion processes.
@@ I Introduction
The existence of low-frequency electro-mechanical oscillations has been recognized as a long-standing industry problem [1], which has recently become even more pronounced. To mention a few, electric power systems have experienced these problems in forms of sub-synchronous resonance (SSR) between turbine modules and series compensated transmission lines [2]; low-frequency oscillations between different control areas during loss of large power plants [3]; and, more recently, wide-area oscillations, such as between Ukraine and Spain, and between Sweden and Spain [4]. The problem of forced oscillations has also been studied [5].
@@
As these problems continue to emerge, it has become very difficult to identify their root causes, and to design control needed to suppress them. Most of the efforts have gone into transient stability simulations and small signal stability analyses, in particular as the slow power plants get replaced by smaller-scale Inverter Based Resources (IBRs) [4]. A particular new challenge perceived by the industry is a lack of detailed models of diverse components. Also, several renown researchers have proposed in the past modeling and analyses approaches, notably use of normal forms [6] and notions of dissipating energy flows [7].
@@
This paper is motivated by the remaining need to have a basic physically-intuitive interpretation of dynamic interactions in terms of energy conversion, power and rate of power exchanges between different energy conversion components and their relations to the existence of inter-area oscillations within an interconnected electric energy system. We attempt to meet this need by observing that each stand-alone module $i$ (component, sub-system) can be characterized in terms of its own internal states and the shared [8] interaction variable $intVar^{i}$ which is a function of its own internal states only, as introduced in our earlier work [9, 10].
@@
In Section II we review a previously introduced aggregate state space model representing dynamics of energy and power within a multi-area interconnected electric power system; we review a previously-introduced structural definition of a a technology-agnostic interaction variable associated with each stand-alone module $i$; and, conceptualize for the first time the notion of an electro-mechanical inter-area oscillation as an interplay between interaction variables of neighboring modules $j\in C_{i}$. We stress that these concepts are general and applicable to systems comprising both conventional generation and renewable resources.
@@
Next, in Section III we illustrate a particular case of linearized power-frequency energy dynamics comprising conventional power plants. This section is a short summary of the earlier work [9] which laid the foundations for claiming the existence of interaction variables in linearized models of electric power grids. An important aspect of this modeling of power dynamics is that the load power deviations from forecast can be explicitly modeled.
@@
Section IV concerns eigenvalue analyses of states contributing to inter-area oscillations. To start with, the existence of an interaction variable associated with a stand-alone module is a direct result of its own power conservation. Important and not surprising, structural existence of an eigenmode in an interconnected system, not present in two disconnected subsystems, is now a direct consequence of its conservation of power, and it is shown mathematically, as well as using simulations. Based on these properties, in sub-section IV-A we propose a method for quickly identifying the modules whose states contribute the most to the inter-area oscillations. This method could be potentially of major use in real-world large-scale systems for which understanding causes of inter-area oscillations becomes a major problem [4].
@@ II Structural property of complex multi-area electric power systems
In electric power systems, such as the one shown in Figure 1, typically single port modules, generators and loads, are interconnected via transmission lines whose real power flows are non linearly dependent on the nodal phase angle differences. All components, except perfect resistors, have state variables and time constants defining their natural and/or forced response given initial conditions. In today’s AC electric power systems these state variables are frequency $\omega(t)$, voltage $V(t)$, real power $P(t)$ and reactive power $Q(t)$. When writing the dynamic model in terms of these variables, one quickly loses physical insights. It therefore becomes quite hard to understand root causes of dynamic problems of interest. To overcome this complexity, we take a modular modeling approach to modeling complex interconnected electric power systems.
@@
We introduce a transformed state space which is derived by combining the dynamics of conventional internal state variables $\tilde{x}_{old}(t)$ in each module $i$ (except the nodal phase angle $\delta_{i}$) and its generated power $P_{i}(t)$. The structural form of this transformed state space is derived by introducing a new, transformed state variable comprising a technology-dependent $\tilde{x}_{old}(t)$ and the power $P_{i}(t)=\frac{dE_{i}(t)}{dt}$
$x_{new}(t)=[\tilde{x}_{old}(t)~{}~{}P(t)]^{\top}$ (1)
@@ II-A General structure of single port dynamic components
Here $E_{i}(t)$, and $P_{i}(t)$ are stored energy in module $i$ and real power generated by module $i$ at the port/node. $P_{in,i}(t)$ is the exogenous power input into module $i$. The term $\frac{Ei(t)}{\tau_{i}}$ stands for Joule losses internal to module $i$, and $\tau_{i}$ is the time constant representing rate at which energy stored in the module dissipates, and can be interpreted as the ratio of stored and dissipated energy [11].
We also recall that the rate of change of energy conversion within each module $\dot{E}_{i}(t)$ is technology-specific and can be expressed in terms of derivatives of its own local state variables $\dot{\tilde{x}}_{old}(t)$. It then follows from the basic conservation law in (4) when combined with energy conversion (2) that the dynamics of the local variables $\tilde{x}_{old}(t)$ can be expressed structurally in terms of its own states and its rate of change of energy, namely power.
$\dot{\tilde{x}}_{old}(t)=\tilde{f}_{i}(\tilde{x}_{old}(t),P_{i}(t),u_{i}(t),m_{i}(t))$ (3)
where $u_{i}(t)$ and $m_{i}(t)$ are the local control and exogenous input, respectively. The key observation when interconnecting modules of diverse energy conversion components is that power and rate of change of power are Tellegen’s quantities, as a direct consequence of power conservation [12]. This means that power output from module $i$ must equal power injected into modules $j\in C_{i}$, where $C_{i}$ is the set of $i$’s neighbouring modules. For the example of electric power grid, power produced by the generator equals the sum of power taken out by the components $j$ directly connected to it, namely transmission lines.
@@
The dynamics of $x_{new}$ takes on the form,
$\dot{x}_{new}(t)$ $=\begin{bmatrix}\dot{\tilde{x}}_{old}(t)\\ \dot{P}_{i}(t)\end{bmatrix}=\begin{bmatrix}\tilde{f}_{i}(\tilde{x}_{old}(t),P_{i}(t),u_{i}(t),m_{i}(t)\\ h_{i}(\tilde{x}_{old}(t),\dot{P}_{j},j\in C_{i})\end{bmatrix}$ (6)
@@ II-B Existence of interaction variable at component level
By re-writing conservation of power at a component level shown in (2) as
@@ II-D Interaction variable of stand-alone area $i$
where $F_{1-2}$ is the power flow from Area 1 to Area 2 through the tie-line. We can infer that, for a disconnected area, $F_{1-2}(t)=constant$ and $\dot{F}_{1-2}(t)=0$, i.e. net power out of disconnected area must remain constant and its derivative is zero. This linear combination of power generated by all modules within a given Area 1 is defined as the $intVar^{1}(t)$ of Area 1 with the rest of the system [9, 10].
@@ II-E Inter-area dynamics: Definition in terms of interaction variables of stand alone areas
We propose, given the above fundamental relations for energy conversion, and the notion of stand-alone of interaction variable $\dot{z}^{I}(t)$, for each area $I$ that when interconnecting the two areas shown in Figure 1 via transmission line whose parameter is $X_{2}$, general Tellegen’s Theorem in (4) implies:
@@ III Linearized modeling of power dynamics
In this section we suggest that the linearized modeling of power dynamics introduced earlier in [9] directly follows from the more general derivations in Section II which do not require linearization and are technology-agnostic. For completeness, we briefly summarize this early derivation. To do so, we consider the same two-area interconnected system shown in Figure 1, which comprises a mix of conventional generators and loads interconnected via lossless transmission lines.
@@ III-B Network constraints
where the network has $n$ generators and $m$ loads. $\delta_{G}$ and $\delta_{L}$ correspond to the vector of phase angles at the generator and load bus respectively. The network constraints given in (13) and (14) are well-known linearized power flow equations, and represent a particular case of a more general power balance given as (4) earlier. The main step in deriving a transformed state space requires applying general constraints on rate of change of power shown in (5) introduced earlier, implying that differentiation of power balance equations is a valid step. In today’s power systems problems $P_{L}$ is predicted and dynamical system response to its deviations $\dot{P}_{L}$ is of direct interest. This is obtained by expressing $\omega_{L}$ as a disturbance caused by the load power deviations $\dot{P}_{L}$ and replacing it in the network constraints imposed on the rate of change of power generation $\dot{P}_{G}$ In short, define
@@ IV-A Methodology to identify the interconnection eigen-mode
  • Disconnected Isolated System (DIS): The same system in Figure 1, but with a disconnected inter-area tie line
@@ V-A Effect of tie-line strength
On doing an eigen-mode and participation factor analysis for case 1, we get the following non-zero eigen-values and their most contributing state variables.
@@ V-B Effect of inertia of generator
The interconnection eigen-mode ($\omega^{2}$) is seen to get slower when $G_{3}$ has a large inertia. Also, $G_{3}$ states start contributing more to the eigen-mode. As is expected, generators with large inertia have a slower response and can be seen in the Figure 3 which plots the inter-area oscillation for the two cases.
@@
VI Impact of Renewable Resources on IAO
The transformed state space has another advantage; it allows us to express load disturbances in the form of $\dot{P}_{L}$ (17). Specifically, renewable sources can be characterised as a varying negative load. Consider the network in Figure 1, with and without a renewable source in the place of $L_{1}$. When the damping is insufficient and the renewable power output deviations include the interconnection mode, e.g. $\dot{P}_{L}=A_{\circ}\sin(\omega t)$ where $\omega$ is the interconnection eigenvalue, it resonates with the IAO and makes it unstable. When the deviations don’t include the interconnection eigen-mode they don’t resonate with the IAO, as can be seen in Figure 4.
@@ VII Concluding remarks
The notion of intVar used in this paper has two unique properties essential for distributed modeling, sensing, computing and control based on local knowledge only: First, when area is disconnected intVar remains constant, and, second, its rate of power exchange with other areas is zero. Given these properties, we claim that the energy dynamics interpretation of interactions within a multi-area system is structural and that it is fundamentally applicable when combining: 1) technology-specific internal dynamics of modules with. 2) generalized Tellegen’s theorem constraints, independent from the types of internal modular energy conversion processes. An example of aggregate mathematical model is shown on a small electric power system comprising two areas to illustrate these structural properties. Further relations between feedback linearizing control for suppressing slow inter-area oscillations between New York and New England areas [3] and for eliminating sub-synchronous resonance [13], and the inter-area structural properties should be studied. For Section V-A, all generators are taken to be identical, i.e., their $A_{LC}$ matrix is the same. The generator inertias are $M=3.2\text{ kg m}^{2}$ and line reactance $X_{1}=6.67\cdot 10^{-2}\;\Omega$. For Section V-B, $X_{1}=X_{2}=6.67\cdot 10^{-2}\Omega$. For case 1, generator inertias are $M=3.2\text{ kg m}^{2}$ and for case 2, $M_{G3}=32\text{ kg m}^{2}$. All simulations show deviations of the quantities around equilibrium. As both the examples consider an undamped system, $D=0$ and the time constants $T_{t},T_{g}$ are very large numbers.

### v2

@@ Abstract
This paper introduces a physically-intuitive notion of inter-area dynamics in systems comprising multiple interconnected energy conversion modules. The idea builds on an earlier general approach of setting their structural properties by modeling internal dynamics in stand-alone modules (components, areas) using the fundamental conservation laws between energy stored and generated, and then constraining explicitly their Tellegen’s quantities (power and rate of change of power). In this paper we derive, by following the same principles, a transformed state-space model for a general nonlinear system. Using this model we show the existence of an area-level interaction variable, $intVar$, whose rate of change depends solely on the area internal power imbalance and is independent of the model complexity used for representing individual module dynamics in the area. Given these structural properties of stand-alone modules, we define in this paper for the first time an inter-area variable as the difference of power wave incident to tie-line from Area I and the power reflected into tie-lie from Area II. Notably, these power waves represent the interaction variables associated with the two respective interconnected areas. We illustrate these notions using a linearized case of two lossless inter-connected areas, and show the existence of a new inter-area mode when the areas get connected. We suggest that lessons learned in this paper open possibilities for computationally-efficient modeling and control of inter-area oscillations, and offer further the basis for modeling and control of dynamics in changing systems comprising faster energy conversion processes.
@@ I Introduction
The existence of low-frequency electro-mechanical oscillations has been recognized and studied as a long-standing industry problem [1, 2], which has recently become even more pronounced. To mention a few, electric power systems have experienced these problems in the form of sub-synchronous resonance (SSR) between turbine modules and series compensated transmission lines [3]; low-frequency oscillations between different control areas during loss of large power plants [4]; and, more recently, wide-area oscillations, such as between Ukraine and Spain, and between Sweden and Spain [5]. The problem of forced oscillations has also been studied [6].
@@
As these problems continue to emerge, it has become very difficult to identify their root causes, and to design control needed to suppress them. Most of the efforts have gone into transient stability simulations and small signal stability analyses, in particular as the slow power plants get replaced by smaller-scale Inverter Based Resources (IBRs) [5]. A particular new challenge perceived by the industry is a lack of detailed models of diverse components. Also, several renown researchers have proposed in the past modeling and analyses approaches, notably use of normal forms [7] and notions of dissipating energy flows [8]. While these methods help with locating the source of inter-area oscillations, they do not model the inter-area interactions dynamically.
@@
The primary contribution of this paper is to meet the need to have a physically-intuitive interpretation of dynamic interactions, in terms of energy conversion, power and rate of power exchanges, between different energy conversion components and their relations to the existence of inter-area oscillations within an interconnected electric energy system. We do this by observing that each stand-alone module $i$ (component, sub-system) can be characterized in terms of its own internal states and the shared [9] interaction variable $intVar^{i}$ which is a function of its own internal states only, as introduced in our earlier work [10, 11]. We then provide a notion of dynamic interaction between two areas by introducing, for the first time, an inter-area variable as the difference of the interaction variables of the respective areas.
@@
In Section II we review a previously introduced aggregate state space model representing dynamics of energy and power within a multi-area interconnected electric power system; we review a previously-introduced structural definition of a technology-agnostic interaction variable associated with each stand-alone module $i$; and, define electro-mechanical inter-area oscillation as an interplay between interaction variables of neighboring modules $j\in C_{i}$. We stress that these concepts are general and applicable to systems comprising both conventional generation and renewable resources.
@@
Next, in Section III, we illustrate a particular case of linearized power-frequency energy dynamics comprising conventional power plants. This section is a short summary of the earlier work [10] which laid the foundations for claiming the existence of interaction variables in linearized models of electric power grids. An important aspect of this modeling of power dynamics is that the load power deviations from forecast can be explicitly modeled.
@@
Section IV concerns eigenvalue analyses of states contributing to inter-area oscillations. To start with, the existence of an interaction variable associated with a stand-alone module is a direct result of its own power conservation. Importantly, the structural existence of an eigenmode in an interconnected system, not present in two disconnected subsystems, is now shown to be a direct consequence of power conservation, both mathematically and through simulations. Based on these properties, in sub-section IV-A we propose a method for quickly identifying the modules whose states contribute the most to the inter-area oscillations. This method could be potentially of major use in real-world large-scale systems for which understanding causes of inter-area oscillations becomes a major problem [5].
@@ II Structural property of complex multi-area electric power systems
In electric power systems, such as the one shown in Figure 1, typically single port modules, generators and loads, are interconnected via transmission lines whose real power flows are nonlinearly dependent on the nodal phase angle differences. All components, except perfect resistors, have state variables and time constants defining their natural and/or forced response given initial conditions. In today’s AC electric power systems these state variables are frequency $\omega(t)$, voltage $V(t)$, real power $P(t)$ and reactive power $Q(t)$. When writing the dynamic model in terms of these variables, one quickly loses physical insights. It therefore becomes quite hard to understand root causes of dynamic problems of interest. To overcome this complexity, we take a modular approach to modeling complex interconnected electric power systems.
@@
We introduce a transformed state space which is derived by combining the dynamics of conventional internal state variables $\tilde{x}_{old,i}(t)$ in each module $i$ (except the nodal phase angle $\delta_{i}$) and its generated power $P_{i}(t)$. The structural form of this transformed state space is derived by introducing a new, transformed state variable comprising a technology-dependent $\tilde{x}_{old,i}(t)$ and the power $P_{i}(t)=\frac{dE_{i}(t)}{dt}$
$x_{new,i}(t)=\begin{bmatrix}\tilde{x}_{old,i}(t)&P_{i}(t)\end{bmatrix}^{\top}$ (1)
@@ II-A General structure of single port dynamic components
Here $E_{i}(t)$, and $P_{i}(t)$ are stored energy in module $i$ and real power generated by module $i$ at the port/node. $P_{in,i}(t)$ is the exogenous power input into module $i$. The term $\frac{E_{i}(t)}{\tau_{i}}$ stands for Joule losses internal to module $i$, and $\tau_{i}$ is the time constant representing rate at which energy stored in the module dissipates, and can be interpreted as the ratio of stored and dissipated energy [12].
We also recall that the rate of change of energy conversion within each module $\dot{E}_{i}(t)$ is technology-specific and can be expressed in terms of derivatives of its own local state variables $\dot{\tilde{x}}_{old,i}(t)$. It then follows from the basic conservation law in (4) when combined with energy conversion (2) that the dynamics of the local variables $\tilde{x}_{old,i}(t)$ can be expressed structurally in terms of its own states and its rate of change of energy, namely power.
$\dot{\tilde{x}}_{old,i}(t)=\tilde{f}_{i}(\tilde{x}_{old,i}(t),P_{i}(t),u_{i}(t),m_{i}(t))$ (3)
where $u_{i}(t)$ and $m_{i}(t)$ are the local control and exogenous inputs, respectively. The key observation when interconnecting modules of diverse energy conversion components is that power and rate of change of power are Tellegen’s quantities, as a direct consequence of power conservation [13]. This means that power output from module $i$ must equal power injected into modules $j\in C_{i}$, where $C_{i}$ is the set of $i$’s neighbouring modules. For the example of electric power grid, power produced by the generator equals the sum of power taken out by the components $j$ directly connected to it, namely transmission lines.
@@
The dynamics of $x_{new,i}$ takes on the form,
$\dot{x}_{new,i}(t)$ $=\begin{bmatrix}\dot{\tilde{x}}_{old,i}(t)\\ \dot{P}_{i}(t)\end{bmatrix}=\begin{bmatrix}\tilde{f}_{i}(\tilde{x}_{old,i}(t),P_{i}(t),u_{i}(t),m_{i}(t)\\ h_{i}(\tilde{x}_{old,i}(t),\dot{P}_{j},j\in C_{i})\end{bmatrix}$ (6)
@@ II-B Existence of interaction variable at component level
Consider a component’s interaction variable as the power generated by the component $P_{i}$ [10, 11]. By re-writing (2) as
@@ II-D Interaction variable of stand-alone area $i$
where $F_{1-2}$ is the power flow from Area 1 to Area 2 through the tie-line. We can infer that, for a disconnected area, $F_{1-2}(t)=constant$ and $\dot{F}_{1-2}(t)=0$, i.e. net power out of disconnected area must remain constant and its derivative is zero. This linear combination of net power generated by all modules within a given Area 1 is defined as the $intVar^{1}(t)$ of Area 1 with the rest of the system [10, 11].
@@ II-E Inter-area dynamics: Definition in terms of interaction variables of stand alone areas
We propose, given the above fundamental relations for energy conversion, and the notion of stand-alone of interaction variable $intVar^{I}(t)$, for each area $I$ that when interconnecting the two areas shown in Figure 1 via transmission line, whose parameter is $X_{2}$, general Tellegen’s Theorem in (4) implies:
@@ III Linearized modeling of power dynamics
In this section we suggest that the linearized modeling of power dynamics introduced earlier in [10] follows from the more general derivations in Section II which do not require linearization and are technology-agnostic. For completeness, we briefly summarize this early derivation. To do so, we consider the same two-area interconnected system shown in Figure 1, which comprises a mix of conventional generators and loads interconnected via lossless transmission lines.
@@ III-B Network constraints
where the network has $n$ generators and $m$ loads. $\delta_{G}$ and $\delta_{L}$ correspond to the vector of phase angles, and $F_{G}$ and $F_{L}$ correspond to the real power flow from neighbouring modules at the generator and load bus respectively. The network constraints given in (13) and (14) are well-known linearized power flow equations, and represent a particular case of a more general power balance given as (4) earlier. The main step in deriving a transformed state space requires applying general constraints on rate of change of power shown in (5) introduced earlier, implying that differentiation of power balance equations is a valid step. In today’s power systems problems $P_{L}$ is predicted and dynamical system response to its deviations $\dot{P}_{L}$ is of direct interest. This is obtained by expressing $\omega_{L}$ as a disturbance caused by the load power deviations $\dot{P}_{L}$ and replacing it in the network constraints imposed on the rate of change of power generation $\dot{P}_{G}$ In short, define
@@ IV-A Methodology to identify the interconnection eigen-mode
  • Disconnected Isolated System (DIS): The same system in Figure 1 but with a disconnected inter-area tie line
@@ V-A Effect of tie-line strength
In our eigen-mode and participation factor analysis for case 1, we identify the following non-zero eigenvalues and their primary contributing state variables.
@@ V-B Effect of inertia of generator
When $G_{3}$ has a large inertia, the interconnection eigen-mode ($\omega^{2}$) gets slower and $G_{3}$ states start contributing more to it. As is expected, generators with large inertia have a slower response and can be seen in the Figure 3 which plots the inter-area oscillation for the two cases.
@@
VI Impact of Intermittent Resources on IAO
The transformed state space has another advantage; it allows us to express load disturbances in the form of $\dot{P}_{L}$ (17). Specifically, renewable sources are intermittent and can be characterised as a varying negative load. Consider the network in Figure 1, with and without a renewable source in the place of $L_{1}$. When the damping is insufficient and the renewable power output deviations include the interconnection mode ($\dot{P}_{L}=A_{\circ}\sin(\omega t)$, $\omega$ is the interconnection eigenvalue), it resonates with the IAO and makes it unstable as can be seen in Figure 4.
@@ VII Concluding remarks
[…]